OpenStudy (linn99123):

What is the approximate circumference of a semicircle with a radius of 35 inches? A. 1923.25 inches B. 219.8 inches C. 109.9 inches D. 54.95 inches

OpenStudy (e.mccormick):

\(C=2\pi r\)

OpenStudy (linn99123):

2*3.14*35?

OpenStudy (anonymous):

B. 219.8

OpenStudy (anonymous):

Remember it is a semi circle (Only 1/2 of a circle)

OpenStudy (e.mccormick):

That is the process Linn.

OpenStudy (e.mccormick):

Ah, I missed that. So yes, the half part is very important!

OpenStudy (linn99123):

it be c then :D

OpenStudy (linn99123):

i feel soo smart :D

OpenStudy (e.mccormick):

That would make stephanie0712's choice very wrong. Circle: \(C=2\pi r\) Semi-Circle (Half-circle): \(C=\dfrac{2\pi r}{2}=\pi r\)

OpenStudy (linn99123):

the replys went in a diffrent order o.o

OpenStudy (e.mccormick):

Yah, the lag does funny things at times. Yes, it is c.